# The Relative Entropy of Expectation and Price

### v4

@@ III Entropic pricing and hedging
Let $d\bar{q}:=dq-qr\,dt$ be the underlying return centralised in the price measure and let $P:=p+dp$ be the final derivative price. Using the price measure, the hedge and price conditions are re-expressed as:
@@
These conditions express a non-linear relationship between the initial and final derivative prices. By definition, the derivative security with notional $\lambda$ has final price $\lambda P$. The corresponding initial price $p_{\lambda}$ and hedge portfolio $\delta_{\lambda}$ per unit notional and the implied funding spread $s_{\lambda}$ are determined from the hedge and price conditions:
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\alpha qs_{\lambda}\,dt)\exp[-\lambda\alpha(P-\delta_{\lambda}\cdot d\bar{q})]]$ (46)
$p_{\lambda}$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\lambda\alpha][P-\delta_{\lambda}\cdot d\bar{q}]$
and the self-funding condition. Scaling the derivative notional is thus equivalent to scaling the risk aversion.
When $\alpha$ is small, the hedge and price conditions are expanded to discover the solutions:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]/(1+r\,dt)-\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\alpha]$
$\delta$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,P]+qs\,dt)+O[\alpha]$ (47)
$p$ $=\frac{1}{1+r\,dt}\left(\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]\right)+O[\alpha^{2}]$
This expansion in terms of $\alpha$ approximates the price when the risk appetite of the investor is large or when the derivative security has small notional. In the marginal market, the model thus implies a limit order book for the derivative security with mid price and bid-offer spread:
@@
The discounted mid price of the derivative security is a martingale in the price measure and the bid-offer spread prices incompleteness of the underlying market. Higher-order adjustments then accommodate the extended features of the price distributions.
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in figures 7, LABEL: and 8 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
III.1 Empty market
In this section, assume that there are no underlying securities available to the investor for funding or hedging.
Assumption III.1 (Empty market).
The market of underlying securities is empty.
It is interesting to note that entropic pricing handles this trivial case in a sensible way. The calibration and hedge conditions are empty, and the price condition implies that the derivative price is a log-martingale in the expectation measure:
@@ III.1 Empty market
In the absence of any market information, price reduces to expectation, adjusted to match the risk appetite of the investor.
III.2 Predictable funding
In this section, assume that the underlying market available to the investor includes a funding security whose final price is predictable.
Assumption III.2 (Predictable funding).
The market partitions into a vector of underlying securities with initial price $q$ and final price $q+dq$ and a funding security with strictly positive initial price $b>0$ and predictable final price $b+db$ that has zero variance $\mathbb{V}[db]=0$ in the expectation measure $\mathbb{E}$.
With this assumption, the calibration and hedge conditions are solved for the implied funding rate $r$ and implied funding spread $s$:
$r$ $=\frac{\mathbb{E}[db]}{b\,dt}$ (50)
$s$ $=0$
When no additional underlying securities are available, the price condition implies that the derivative price is a discounted log-martingale in the expectation measure:
$p=\frac{1}{1+r\,dt}\mathbb{E}[\alpha][P]$ (51)
More generally, underlying securities are available for hedging, and the calibration and hedge conditions for the unit optimal portfolio $\phi$ and hedge portfolio $\delta$ become:
00 0 $=\mathbb{E}^{\phi}[d\bar{q}]$ (52)
00 0 $=\mathbb{E}^{\phi}[d\bar{q}\,\exp[-\alpha(P-\delta\cdot d\bar{q})]]$
The price condition then derives the discounted log-martingale expression in the price measure:
$p=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\alpha][P-\delta\cdot d\bar{q}]$ (53)
The funding security funds the investment and manifests as discounting in the price expression, and the additional underlying securities hedge the derivative risks.
III.3 Continuous diffusion
In this section, assume that the underlying and derivative prices diffuse continuously, and the investor is able to re-balance the hedge over short investment horizons.
Assumption III.3 (Continuous diffusion).
The diffusion of the underlying and derivative prices is continuous almost certainly in the expectation measure $\mathbb{E}$, so that their cumulant generating function satisfies:
@@ III.3 Continuous diffusion
with $\mathbb{V}[dp]=O[dt]$, $\mathbb{V}[dq,dp]=O[dt]$ and $\mathbb{V}[dq]=O[dt]$.
Two assumptions are included here. The first assumption, that the market prices diffuse continuously, justifies the short-horizon expression for their cumulant generating function. The second assumption is that the investor re-balances frequently and at negligible cost, so that the $O[dt]$ corrections to the investment strategy can be ignored, and the bid-offer spread around the mid price derives exclusively from the post-hedging residual risks.
@@
These results are sufficient to solve the conditions. The calibration condition is solved for the implied funding rate $r$ and the unit optimal portfolio $\phi$:
@@
and the hedge and price conditions are solved for the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$ (58)
$\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)+O[dt]$
@@
III.4 Markovian diffusion
In this section, assume that the underlying and derivative prices are deterministic functions of a state vector.
Assumption III.4 (Markovian diffusion).
There is a vector $x$ of state variables whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
@@ III.4 Markovian diffusion
where the vector $\mu\equiv\mu[t,x]$ and positive-definite matrix $\nu\equiv\nu[t,x]$ are the mean and covariance rate for the continuous component of the diffusion, and the positive measure $\theta[dj]\equiv\theta[t,x][dj]$ is the frequency of jumps in the discontinuous component of the diffusion. The underlying price $q\equiv q[t,x]$ and derivative price $p\equiv p[t,x]$ are assumed to be functions of the time $t$ and state $x$.
This form for the cumulant generating function of the state implies its distribution is divisible, which enables the transition to the limit $dt\to 0$ of continuous hedging. Repeated differentiation with respect to $\eta$ generates Itô’s lemma for the function $f\equiv f[t,x]$ of the time $t$ and state $x$ in the expectation measure $\mathbb{E}$:
@@
$0={}$ $\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+qs-\left(\frac{\partial q}{\partial x}\cdot\nu\frac{\partial q}{\partial x}\right)\delta$ (64)
@@
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}p}{\partial x^{2}}\cdot{}\!\cdot\nu$
@@
which are solved simultaneously with the self-funding condition $\delta\cdot q=p$ for the implied funding spread $s$, hedge portfolio $\delta$ and derivative price $p$.
@@
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a linear parabolic differential-integral equation for the derivative price. This equation balances the contributions to the return from the continuous and discontinuous components of the diffusion. The underlying market is incomplete and the hedge strategy does not eliminate market risk, but this choice for the hedge portfolio optimises the return after hedging.
@@
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a non-linear parabolic differential equation for the derivative price. This equation modifies the expression for the price to adjust for residual risk. The underlying market is incomplete and the hedge strategy does not eliminate market risk, but this choice for the hedge portfolio optimises the return after hedging.
III.5 Heston model
In this section, the Heston model is proposed for the expectation measure. This concrete example of a continuous Markovian diffusion illustrates the key features of the more general framework.
Assumption III.5 (Heston model).
The state comprises a funding price $b$, an underlying price $q$ and a variance $\nu$ whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
@@ III.5 Heston model
for the hedge portfolio $\delta$ and the derivative price $p$. In the first expression, the hedge ratio is adjusted when the increment of the variance is correlated with the increment of the underlying price. The derivative price then satisfies a linear parabolic differential equation with a non-linear correction term that prices the unhedgeable risks when the increments are not perfectly correlated.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 1,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@
Figure 11: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@
Figure 12: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.

### v5

@@ III.1 Calibration
If the underlying market includes a risk-free funding security with initial price $b>0$ and final price $B=b+db$ satisfying $\mathbb{V}[db]=0$, then the calibration condition derives the solution $r=\mathbb{E}[db]/(b\,dt)$ for the implied funding rate. More generally, the implied funding rate is discovered simultaneously with the unit optimal portfolio by iterating the Newton-Raphson scheme:
$\partial r$ $=\frac{\displaystyle(\phi+\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-qr\,dt))\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}$ (48)
$\partial\phi$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-q(r+\partial r)\,dt)$
initialised with $r=0$ and $\phi=0$.
As evident in this calibration scheme, the implied funding rate incorporates contributions from all the underlying returns, weighted by their predictabilities so that the least risky returns dominate. This property is demonstrated in the studies of figure 6, which simplify the model with the assumption that the final underlying prices are independent. The calibration conditions decouple to generate a set of independent conditions:
$r=\frac{1}{dt}\left(\frac{\displaystyle\mathbb{E}[Q_{i}\exp[-\phi_{i}Q_{i}]]}{\displaystyle q_{i}\mathbb{E}[\exp[-\phi_{i}Q_{i}]]}-1\right)$ (49)
that are solved separately to express the components of the unit optimal portfolio $\phi[r]$ as functions of the implied funding rate $r$. The self-funding condition $\phi[r]\cdot q=0$ is then solved for $r$.
In the first study, the final underlying price $Q_{i}$ is modelled as a normal variable with mean $q_{i}(1+\mu_{i}\,dt)$ and variance $q_{i}^{2}\sigma_{i}^{2}\,dt$, where $\mu_{i}$ is the drift and $\sigma_{i}$ is the volatility. The calibration condition derives the unit optimal portfolio:
$\phi_{i}=\frac{\mu_{i}-r}{q_{i}^{\vphantom{2}}\sigma_{i}^{2}}$ (50)
and the self-funding condition then derives the implied funding rate:
$r=\sum\nolimits_{i}\frac{\displaystyle 1/\sigma_{i}^{2}}{\displaystyle\sum\nolimits_{j}1/\sigma_{j}^{2}}\,\mu_{i}$ (51)
The implied funding rate is the weighted average of the drifts, with weight inversely proportional to the square of the volatility.
In the second study, the final underlying price $Q_{i}$ is modelled as a binomial variable with possible values $(0,1+c_{i}\,dt)$ occurring with probabilities $(w_{i},1-w_{i})$, where $c_{i}$ is the coupon and $w_{i}$ is the default probability. The calibration condition derives the unit optimal portfolio:
$\phi_{i}=\frac{1}{1+c_{i}\,dt}\log\!\left[\frac{1-w_{i}}{w_{i}}\left(\frac{1+c_{i}\,dt}{q_{i}(1+r\,dt)}-1\right)\right]$ (52)
and the self-funding condition:
$0=\sum\nolimits_{i}\frac{q_{i}}{1+c_{i}\,dt}\log\!\left[\frac{1-w_{i}}{w_{i}}\left(\frac{1+c_{i}\,dt}{q_{i}(1+r\,dt)}-1\right)\right]$ (53)
is solved for the implied funding rate.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Call Option (Normal) Call Option (Poisson) Call Option (Uniform)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Digital Option (Normal) Digital Option (Poisson) Digital Option (Uniform)
Figure 7: Three studies of entropic hedging and pricing for a call option (top three rows) with final price $P=(Q-k)^{+}$ and a digital option (bottom two rows) with final price $P=1_{Q>k}$ at strike $k$ on an underlying security with final price $Q$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable, all standardised by linear transformation to have zero mean and unit standard deviation. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer around the mid price and skews the optimal hedge ratio for positions with large notional.
III.2 Pricing and hedging
Let $d\bar{q}:=dq-qr\,dt$ be the underlying return centralised in the price measure. Using the price measure, the hedge and price conditions are re-expressed as:
@@ III.2 Pricing and hedging
These conditions express a non-linear relationship between the initial and final derivative prices. By definition, the derivative security with notional $\lambda$ has final price $\lambda P$. The corresponding initial price $p$, hedge portfolio $\delta$ and implied funding spread $s$ per unit notional are determined from the hedge and price conditions:
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\lambda\alpha qs\,dt)\exp[-\lambda\alpha(P-\delta\cdot d\bar{q})]]$ (55)
$p$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\lambda\alpha][P-\delta\cdot d\bar{q}]$
and the self-funding condition. Scaling the derivative notional is thus equivalent to scaling the risk aversion. This is the scale invariance property of entropic pricing.
When $\lambda\alpha$ is small, the hedge and price conditions are expanded to discover the solutions:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]/(1+r\,dt)-\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\lambda\alpha]$
$\delta$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,P]+qs\,dt)+O[\lambda\alpha]$ (56)
$p$ $=\frac{1}{1+r\,dt}\left(\mathbb{E}^{\phi}[P]-\frac{1}{2}\lambda\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]\right)+O[(\lambda\alpha)^{2}]$
This expansion in terms of $\lambda\alpha$ approximates the price when the risk appetite of the investor is large or when the derivative security has small notional. In the marginal market, the model thus implies a limit order book for the derivative security with mid price and bid-offer:
@@
The discounted mid price of the derivative security is a martingale in the price measure and the bid-offer prices incompleteness of the underlying market. Higher-order adjustments then accommodate the extended features of the price distributions.
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The studies in figure 7 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer around the mid model that prices the residual risk after hedging.
It is interesting to note that entropic pricing handles the trivial case of an empty underlying market in a sensible way. The calibration and hedge conditions are empty, and the price condition implies that the derivative price is a log-martingale in the expectation measure:
@@
In the absence of any market information, price reduces to expectation adjusted to match the risk appetite of the investor.
When the underlying market is not empty, the hedge and price conditions may not have simple solutions and are instead solved numerically via the Newton-Raphson scheme. The conditions can be solved explicitly when the prices diffuse continuously, however, and this case is useful to develop intuition. The diffusion of the underlying and derivative prices is continuous in the expectation measure $\mathbb{E}$ when their cumulant generating function satisfies:
@@
with $\mathbb{V}[dp]=O[dt]$, $\mathbb{V}[dq,dp]=O[dt]$ and $\mathbb{V}[dq]=O[dt]$. If the investor re-balances frequently and at negligible cost, the $O[dt]$ corrections to the investment strategy can be ignored, and the bid-offer around the mid price derives exclusively from the post-hedging residual risks.
@@
These results are sufficient to solve the conditions. The calibration condition is solved:
@@
for the implied funding rate $r$ and the unit optimal portfolio $\phi$, and the hedge and price conditions are solved:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$
$\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)+O[dt]$ (63)
@@
for the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$.
III.3 Market and model risk
Starting with the expectation measure $\mathbb{E}$, the calibration conditions:
00 0 $=\phi\cdot q$ (64)
00 0 $=\mathbb{E}^{\phi}[dq]-qr\,dt$
$p$ $=\delta\cdot q$ (65)
$p$ $=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$
are solved to generate the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$ for the derivative security. While the hedge neutralises the underlying component of the derivative risk, when the market is incomplete there are unhedged risks that are sensitive to the choice of expectation measure. The entropic pricing framework distinguishes between these market and model risks in the derivative security.
$\partial\phi$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,W]-q\,\partial r\,dt)+O[\epsilon]$
$\bar{\mathbb{E}}^{\bar{\phi}}[X]-\mathbb{E}^{\phi}[X]=$ (68)
$\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][X]-\mathbb{E}^{\phi}[\alpha][X]=$
for the indeterminate variable $X$.
$\bar{p}$ $=\bar{\delta}\cdot q$ (69)
$\bar{p}$ $=\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][P-\bar{\delta}\cdot dq]$
$\partial p=\frac{1}{1+(r+\alpha s)\,dt}\times$ (70)
IV Markovian diffusion
In this section, assume that there is a vector $x$ of state variables whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
@@ IV Markovian diffusion
where the vector $\mu\equiv\mu[t,x]$ and positive-definite matrix $\nu\equiv\nu[t,x]$ are the mean and covariance rate for the continuous component of the diffusion, and the positive measure $\theta[dj]\equiv\theta[t,x][dj]$ is the frequency of jumps in the discontinuous component of the diffusion. The underlying price $q\equiv q[t,x]$ and derivative price $p\equiv p[t,x]$ are then assumed to be functions of time $t$ and state $x$.
This Lévy-Khintchine form for the cumulant generating function of the state implies its distribution is divisible, which enables the transition to the limit $dt\to 0$ of continuous hedging. Repeated differentiation with respect to $\eta$ derives Itô’s lemma for the function $f\equiv f[t,x]$ of time $t$ and state $x$ in the expectation measure $\mathbb{E}$:
@@
$0={}$ $\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+qs-\left(\frac{\partial q}{\partial x}\cdot\nu\frac{\partial q}{\partial x}\right)\delta$
@@
which are solved simultaneously with the self-funding condition $\delta\cdot q=p$ for the implied funding spread $s$, hedge portfolio $\delta$ and derivative price $p$. Incompleteness means that the hedge portfolio does not exactly offset the risks of the derivative security for all possible market movements over the hedging interval, and this residual risk is accounted for with a compensating adjustment to the derivative price.
@@
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a linear parabolic differential-integral equation for the derivative price. This equation balances the contributions to the return from the continuous and discontinuous components of the diffusion.
@@
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a non-linear parabolic differential equation for the derivative price. This equation modifies the expression for the price to adjust for residual risk.
As a concrete example of a continuous Markovian diffusion, the Heston model illustrates the key features of the more general framework. In this model, the state comprises a funding price $b$, an underlying price $q$ and a variance $\nu$ satisfying the stochastic differential equations:
$db$ $=br\,dt$ (81)
$dq$ $=q(\mu\,dt+\sqrt{\nu}\,dy)$
$d\nu$ $=(\bar{\nu}-\nu)\kappa\,dt+\sqrt{\nu}\,\sigma\,dz$
where $dy$ and $dz$ are Brownian increments in the expectation measure with correlation $\rho$. This is equivalent to the cumulant generating function:
@@
for the hedge portfolio $\delta[t,b,q,\nu]$ and the derivative price $p[t,b,q,\nu]$. In the first expression, the hedge ratio is adjusted when the increment of the variance is correlated with the increment of the underlying price. The derivative price then satisfies a linear parabolic differential equation with a non-linear correction term that prices the unhedgeable risks when the increments are not perfectly correlated.
V Discrete models
As the example of the Heston model demonstrates, entropic risk optimisation can be consistently applied in theoretical models that include unhedgeable factors, naturally adjusting for market incompleteness in pricing and hedging. Data-driven models take a different approach, eschewing theoretical foundations for a statistical method that extracts information directly and exclusively from the data. In practice, this means that the model is supported on a finite set of eigenstates generated from a statistical model trained on historical data. The expectation measure is then assumed to be uniformly distributed on the eigenstates. With no theory to support development beyond the information expressed in the data, a more general foundation is needed that does not depend on unrealistic assumptions such as complete markets, continuous hedging or risk-free funding. Advocates of deep hedging address this by replacing risk neutralisation with risk optimisation, using convex risk metrics to assess strategic performance. While they suffer from lack of explainability, data-driven models can uncover complex relationships between economic variables that may be hard to capture theoretically.
@@ V Discrete models
Distribution (100 states) Distribution (1,000 states) Distribution (10,000 states)
Figure 8: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=100$, $1,000$ and $10,000$ samples from a standardised normal variable, centralised to have zero mean. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=-0.4$ (top row), $q=0$ (middle row) and $q=0.4$ (bottom row). In the transformation from expectation to price measure, the uniform probabilities are biased using an exponential function of the final underlying price. Among equivalent measures that satisfy the calibration condition, the price measure thus has minimum entropy relative to the expectation measure.
@@
Call Option (100 states) Call Option (1,000 states) Call Option (10,000 states)
@@
Digital Option (100 states) Digital Option (1,000 states) Digital Option (10,000 states)
Figure 9: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=100$, $1,000$ and $10,000$ samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=-0.4$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option (100 states) Call Option (1,000 states) Call Option (10,000 states)
@@
Digital Option (100 states) Digital Option (1,000 states) Digital Option (10,000 states)
Figure 10: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=100$, $1,000$ and $10,000$ samples from a standardised normal variable. The price measure is supported on the same samples with uniform probabilities that match the target initial underlying price $q=0$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option (100 states) Call Option (1,000 states) Call Option (10,000 states)
@@
Digital Option (100 states) Digital Option (1,000 states) Digital Option (10,000 states)
Figure 11: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=100$, $1,000$ and $10,000$ samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=0.4$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.